# Uncertainty in Security: managing cyber senescence

### v1

@@ 2. Security
The NIS2 and DORA regulations also contain clauses that make executive leaders in an organisation personally liable for blameworthy failing security setups. Hence the lawmaker is forcing the cybersecurity topic to be on the agenda of the executive board. Appearantly, this required the force of law. This brings me to the central theme of my lecture: how are humans supposed to make good security decisions?
@@
In that year, the mathematician Kurt Gödel proved his famous incompleteness theorem. This theorem states that any consistent formal system, of sufficient strength or complexity [^2] This is an inprecise formulation. The theorem assumes the formal system is capable of expressing propositions about the natural numbers., contains true statements that cannot be proved in that system. If even mathematics is not decidable or calculable, how can one hope to reduce human reasoning to calculations?
@@
Can systems be secure and in use? In 2015, the researcher Cormac Herley published an elegant paper ([herley]) in the Proceedings of the National Academy of Sciences in the USA. In that paper he shows that a claim that states that a system is not secure, cannot be falsified. He showed that the set $Y$ is undecidable, where $Y$ is defined as the set of system statusses $x$ of a system $X$, with the property
@@
Executives that are personally liable, will continue to have to deal with high levels of uncertainty in cybersecurity decision making, and they will rather be safe than sorry. As we have seen, this means there is a natural selection pressure on the cybersecurity system to add controls, while falsifiability of security claims adds friction to removal of controls. This mechanism will lead to an accumulation of waste in the system: controls that are no longer necessary and could be replaced or removed but will stay.
@@
How is life of an octopus a metaphor of cyber security? Cyber security has become a revenue model for the IT industry. Uncertainty about the real risks and the effectiveness of controls, combined with risk averseness of executives who start to feel their personal liability leads to high selection pressure to quickly patch vulnerabilities for every newly discovered threat. The agency problems that arise in security decision making and the biases that play into this are subject of recent study, see for example [elif] and [barre]. The falsifiability of security claims results in only adding controls regardless their long-term downsides. This leads to an accumulation of waste in the cyber security system. In other words: the cyber security system evolves in an eco-system that looks a lot like the environment in which an octopus evolved and in which all incentives point in the wrong direction.
@@
Firstly, we need to look for ways to make better local security decisions under uncertainty. This means we need to become better at detection of incidents, for example by applying AI to analyse the increasing volumes of telemetry, and make better threat analysises to reduce uncertainty and help choose controls. A lot of research is focussing on these problems and recently Lampis Alevizos and myself have published papers about this ([dekker] and [alevizos]).

### v2

@@ 2. Security
The NIS2 and DORA regulations also contain clauses that make executive leaders in an organisation personally liable for blameworthy failing security setups. Hence the lawmaker is forcing the cybersecurity topic to be on the agenda of the executive board. Apparantly, this required the force of law. This brings me to the central theme of my lecture: how are humans supposed to make good security decisions?
@@
In that year, the mathematician Kurt Gödel proved his famous incompleteness theorem. This theorem states that any consistent formal system, of sufficient strength or complexity [^2] This is an imprecise formulation. The theorem assumes the formal system is capable of expressing propositions about the natural numbers., contains true statements that cannot be proved in that system. If even mathematics is not decidable or calculable, how can one hope to reduce human reasoning to calculations?
@@
Can systems be secure and in use? In 2015, the researcher Cormac Herley published an elegant paper ([herley]) in the Proceedings of the National Academy of Sciences in the USA. In that paper he shows that a claim that states that a system is not secure, cannot be falsified. He showed that the set $Y$ is undecidable, where $Y$ is defined as the set of system statuses $x$ of a system $X$, with the property
@@
Executives that are personally liable, will continue to have to deal with high levels of uncertainty in cybersecurity decision making, and they will rather be safe than sorry. As we have seen, this means there is a natural selection pressure on the cybersecurity system to add controls, while unfalsifiability of security claims adds friction to removal of controls. This mechanism will lead to an accumulation of waste in the system: controls that are no longer necessary and could be replaced or removed but will stay.
@@
How is life of an octopus a metaphor of cyber security? Cyber security has become a revenue model for the IT industry. Uncertainty about the real risks and the effectiveness of controls, combined with risk averseness of executives who start to feel their personal liability leads to high selection pressure to quickly patch vulnerabilities for every newly discovered threat. The agency problems that arise in security decision making and the biases that play into this are subject of recent study, see for example [elif] and [barre]. The unfalsifiability of security claims results in only adding controls regardless their long-term downsides. This leads to an accumulation of waste in the cyber security system. In other words: the cyber security system evolves in an eco-system that looks a lot like the environment in which an octopus evolved and in which all incentives point in the wrong direction.
@@
Firstly, we need to look for ways to make better local security decisions under uncertainty. This means we need to become better at detection of incidents, for example by applying AI to analyse the increasing volumes of telemetry, and make better threat analyses to reduce uncertainty and help choose controls. A lot of research is focussing on these problems and recently Lampis Alevizos and myself have published papers about this ([dekker] and [alevizos]).